OpenStudy (anonymous):

jimmy's father's age is 6 years more than twice as jimmy's age. The product of their age is 360. How old is Jimmy?

OpenStudy (amistre64):

f = 2j + 6 fj = 360

OpenStudy (amistre64):

solve for f in terms of j and sub it into the other equation

OpenStudy (anonymous):

Let jimmy's age be x. So, his father's age = 2x + 6 x(2x + 6) = 360 2x^2 + 6x = 360 2x^2 + 6x - 360 = 0 Now, you can solve it using the quadratic formula.

OpenStudy (anonymous):

Let Jimmy's age be x Jimmy's father's age is x * 2 + 6 Therefore, x * x * 2 + 6 = 360 x * x * 2 = 360 - 6 2x * 2 = 354 x = (and now u can solve it)

OpenStudy (amistre64):

one thing I never understood was why we would use a variable that is not indicative of the subjects of the problem :)

OpenStudy (amistre64):

gary, that is not a product of the two ages

OpenStudy (amistre64):

you have the +6 just being a spurious little klingon

OpenStudy (anonymous):

yes, you have made a mistake.

OpenStudy (anonymous):

ooooh yup i hv made a mistake...

OpenStudy (anonymous):

When doing word problems like this there are key words to look for. First we know we need two varables because we need Jimmy's father (F) and Jimmy ( J ) Example of breaking the sentence down jimmy's father's = f is = equal sign 6 year more = +6 twice as jimmy's age = 2j product of = Means to times The product of their age = f x j = 360

OpenStudy (amistre64):

Doodle, good job, but watch those variable; its bad form to define J and then use j in its place.

OpenStudy (anonymous):

let jimmy's age be x then his father's age is 2x+6 according to given condition x(2x+6)=360 2x^2+6x=360 2(x^2+3x)=360 x^2+3x=180 x^2+3x-180=0 solve further by using quadratic formula